OpenStudy (anonymous):

4y^2-3{2y^2-[(4y^2 - y + 2) - (3y^2 + 2y - 5)]}

OpenStudy (anonymous):

okay imm stuck on a part

OpenStudy (anonymous):

Heres my work : 4y^2 (- 6y2) (-12y^2 - 3y + 6) (-9y^2 + -6y) (-15) (4y^2 - 6y^2 - 12y^2 - 9y^2) (-3+-6y) (6 - (-15)

OpenStudy (anonymous):

this is tricky

OpenStudy (anonymous):

(-3 + 6y) is supposedd to be (-3y + -6y) I collected the like terms, but i think im doing something wrong. the answer is supposed to be y^2 - 9y +21

OpenStudy (anonymous):

so whats the objest we have to get?

OpenStudy (anonymous):

just simplifying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(16y-36y-144y-81y)(-9y)(21)

OpenStudy (anonymous):

k got it almost simplied

OpenStudy (anonymous):

I have to multiply it? lol ahh omg i was trying to add / subtract

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is it because of the y^2?

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

the answer sheet says its supposed to be y^2 - 9y + 21

OpenStudy (anonymous):

well i have the 9y+21 part

OpenStudy (anonymous):

what i did was subtract 4y2 - 6y2 and it gave me 2 and then 12y2 - 9y2 and it gave me 3 and i subtracted that and it gave me -y2 but it doesnt sound right

OpenStudy (anonymous):

because the y2 is supposed to be positive

OpenStudy (anonymous):

no no thats right

OpenStudy (anonymous):

it turns postive

OpenStudy (anonymous):

oh whaat the heck lol...

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

so u should have y^2-9y+21

OpenStudy (anonymous):

thx for medal

OpenStudy (anonymous):

yup.. but okay if i subtracted 12 - 9 its 3 right and 4-6 -2? -2 - 3 = 5?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i gtg play some destiny on my ps4

OpenStudy (anonymous):

peace

OpenStudy (anonymous):

i have a question about getting y^2...

OpenStudy (anonymous):

i can help you with that awser

OpenStudy (anonymous):

what i did was (-2y^2 + 3y^2) and got y^2 is that correct?

OpenStudy (anonymous):

but i dont get it because i worked it out and it said 4y^2 - 6y^2 -(-12y^2 - 3y + 6) - (9y^2 - (-15)

OpenStudy (anonymous):

4y2−3(2y2−(4y2−y+2−(3y2+2y−5))) Distribute: =4y2+(−3)(2y2)+(−3)(−4y2)+(−3)(y)+(−3)(−2)+(−3)(3y2)+(−3)(2y)+(−3)(−5) =4y2+−6y2+12y2+−3y+6+−9y2+−6y+15 Combine Like Terms: =4y2+−6y2+12y2+−3y+6+−9y2+−6y+15 =(4y2+−6y2+12y2+−9y2)+(−3y+−6y)+(6+15) =y2+−9y+21 Answer: =y2−9y+21

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

yup